OpenStudy (anonymous):

In a 500-pound mixture of ferric oxide, the ratio of iron to oxygen is

OpenStudy (anonymous):

\[2 Fe _{2}O _{3}=4 Fe+3O _{2}\] 2[2*32+3*16]=4*32+6*16 2(64+48)=128+96 2*112=128+96 224=128+96 iron:oxygen::128:96 or 4:3